# DaMO: A Data-Efficient Multimodal Orchestrator for Temporal Reasoning with Video LLMs

### v1

@@ Abstract
Large Language Models (LLMs) have recently been extended to the video domain, enabling sophisticated video-language understanding. However, existing Video LLMs often exhibit limitations in fine-grained temporal reasoning, restricting their ability to precisely attribute responses to specific video moments, especially under constrained supervision. We introduce DaMO, a data-efficient Video LLM explicitly designed for accurate temporal reasoning and multimodal understanding. At its core, the proposed Temporal-aware Fuseformer employs a hierarchical dual-stream architecture that progressively captures temporal dynamics within each modality and effectively fuses complementary visual and audio information. To further enhance computational efficiency, DaMO integrates a global residual that reduces spatial redundancy while preserving essential semantic details. We train DaMO via a structured four-stage progressive training paradigm, incrementally equipping the model with multimodal alignment, semantic grounding, and temporal reasoning capabilities. This work also contributes multiple datasets augmented from existing ones with GPT-generated temporally grounded QA pairs for tasks requiring temporal supervision. Comprehensive experiments on temporal grounding and video QA benchmarks demonstrate that DaMO consistently surpasses prior methods, particularly in tasks demanding precise temporal alignment and reasoning. Our work establishes a promising direction for data-efficient video-language modeling. Our project is released at https://github.com/yue-723/DaMO
Figure 1: A comparison of temporal reasoning in video-grounded QA. Given a video clip and a temporal-related question, DaMO produces a more accurate and temporally aligned response than Video-LLaMA [56] and VTimeLLM [14], demonstrating its superior temporal understanding.
@@ 1 Introduction
To make these trainings possible, this work also contributes multiple temporal QA datasets by enriching existing datasets through GPT-based prompting, serving as a valuable asset for temporal reasoning benchmarking.
Extensive experiments on temporal grounding and video-dialogue benchmarks validate DaMO’s strong performance, especially in tasks requiring precise temporal reasoning under limited data and computational resources. Fig. 1 demonstrates DaMO’s superior ability to understand and reason over temporal structures in a video.
@@
  • We devise a global residual to preserve global context while significantly reducing computational cost.
  • We present a four-stage progressive training paradigm to systematically enhance DaMO’s multimodal alignment and temporal reasoning capabilities. The training is facilitated by several newly curated datasets with GPT-generated temporal QA pairs.
@@ 3 Multimodal Orchestrator DaMO
DaMO is designed specifically for data-efficient training and effective temporal reasoning for Video LLM. At its core, DaMO features multimodal T-Fuseformer, a dual-stream transformer architecture for robust temporal modeling and multimodal fusion combined with an efficient global residual strategy for spatial dimension reduction. Furthermore, DaMO incorporates a four-stage progressive training strategy, achieving state-of-the-art results with significantly fewer training resources.
Fig. 2 illustrates the architecture of DaMO. The input is an audio-visual stream and a user prompt. A pretrained LLM is incorporated to provide a response. To enable the LLM to reason about the audio-visual stream, we first separate it into a silent video (or simply video) stream and an audio stream. Two pretrained visual and audio encoders are employed to separately extract their modality-specific features. These extracted features undergo a spatial path and a temporal path, and then are processed by Q-Former to project multimodal features into tokens that can adapt to the embedding space of the pretrained LLM accompanied by LoRA.
@@ 3.1 Pretrained Video and Audio Encoders
Each frame $V_{i},i=1..N,$ and audio segment $A_{j},j=1..M,$ is independently encoded by its corresponding encoder. In our implementation, each visual frame is resized to $336\times 336$, and each audio segment length $S=30$ seconds with $R=16$ KHz. The resulting visual and audio features are denoted respectively as:
@@ 3.2 Spatial Pathway
To effectively compress spatial redundancy, we apply adaptive average pooling to $V^{\text{Local}}$ and $A^{\text{Local}}$ separately, reducing both their lengths from $L_{v}-1$ and $L_{a}$ to the same smaller $L^{\prime}$. Simultaneously, the global features $V^{\text{Global}}$ and $A^{\text{Global}}$ are each refined through a lightweight feed-forward network (FFN). Finally, we fuse the compressed local representations with the refined global contexts through a residual addition:
@@
Here, $X=V$ or $A$, $V^{\text{Res}}\in\mathbb{R}^{N\times L^{\prime}\times D_{v}}$, and $A^{\text{Res}}\in\mathbb{R}^{M\times L^{\prime}\times D_{a}}$.
@@ 3.4 Token Projection and LoRA Adaptation
The FUSION queries from the final T-Fuseformer layer thus serves as a compact, temporally grounded representation of the entire video, which is subsequently projected into the LLM’s semantic space using the Q-Former, followed by a linear transformation. These projected multimodal embeddings serve as soft prompts, which are concatenated with user-generated textual queries and fed into the pretrained LLM. To adapt to these FUSION queries, the frozen LLM is enhanced by LoRA [13] to generate temporally grounded and contextually precise responses.
@@ 3.5 Data-Efficient Progressive Training Paradigm
We employ the VTG objective to condition the frozen LLM, guiding it to generate temporally coherent and contextually accurate responses. To facilitate this semantic alignment, we curate a training corpus comprising approximately 300K QA pairs. This dataset integrates several sources, including general video-based QA data from VideoInstruct-100K [37] and a refined version of AVSD processed by Macaw-LLM [35]. To enrich the temporal grounding component, we additionally incorporate four temporally annotated datasets: QVHighlight [19], Charades-STA [12], ActivityNet [18], and 100K samples selected from Koala36M [50]. Building upon these resources, we leverage GPT-based prompting to regenerate approximately 150K QA pairs with explicit temporal grounding derived from their original segment annotations. The augmentation details are provided in Supplementary.
By combining general QA datasets with the GPT-augmented datasets, we equip DaMO with diverse semantic understanding and fine-grained temporal alignment capabilities, providing robust representation bridging.
@@
The objective in this stage is to adapt DaMO for temporally grounded multi-turn dialogue, enabling it to produce coherent responses while reasoning over temporally structured video features. Building upon the training outcomes of Stage 3, we continue fine-tuning the full pipeline using the VTG objective. This stage leverages a curated 39K dialogue-oriented dataset, comprising general dialogues from VideoChat2 [24] and ActivityNet-based dialogues adapted from VTimeLLM [14]. Additionally, we augment DSTC10-AVSD [44] with GPT-based prompting to inject temporal annotations aligned with its dialogue flows, ensuring it to support training temporally grounded dialogue generation. The details are provided in Supplementary.
@@ 4.2 Zero-Shot Video Retrieval Benchmarks
We first evaluate DaMO’s capability to extract discriminative multimodal representations. This evaluation specifically assesses the effectiveness of stage-1 training, focusing on DaMO’s multimodal temporal alignment ability without any retrieval-specific fine-tuning. We test the zero-shot video retrieval task on MSR-VTT [53] and MSVD [3]. Following standard protocols, we report Recall@1, Recall@5, and Recall@10 metrics.
@@
As shown in Table 1, DaMO achieves competitive results despite training on considerably smaller data. On MSR-VTT, DaMO ranks second only to InternVideo2-6B, while on MSVD, it surpasses InternVideo2-6B at Recall@1 and Recall@5, highlighting its strong generalization ability and data efficiency. These results validate the effectiveness of our temporal fusion architecture and multimodal alignment strategy under resource-constrained conditions.
@@ 4.3 Video-LLM Temporal Grounding Benchmarks
As shown in Table 2, DaMO achieves state-of-the-art results on both benchmarks, significantly outperforming existing Video LLM methods. These outcomes affirm our approach’s superior temporal localization performance, validating the effectiveness of our temporal fusion mechanism and progressive training paradigm.
4.4 Video Dialogue Benchmarks
@@ 5 Conclusions
Additionally, this work contributes several temporal QA datasets that are enhanced from existing video-language datasets via GPT-based prompting. Further dataset information is available in Supplementary. In the future, we plan to further enhance the semantic comprehension and reasoning capabilities of DaMO beyond temporal aspects and incorporate additional modalities.
@@ 6 Additional Qualitative Examples
7 GPT-augmented Temporal QA Datasets
To train DaMO to learn to reason about temporal information, we need to curate a temporal QA dataset without depending on costly large-scale manual annotations. There exist several event-level temporally annotated datasets. While these datasets contain valuable temporal signals, their original annotations are either limited to span-based descriptions without QA structure, or lack fine-grained temporal grounding within dialogues. This restricts their direct applicability to temporal QA tasks, especially for those requiring alignment between natural language queries and specific video moments. Therefore, we construct a GPT-augmented dataset from the following datasets: (i) QVHighlight [19], (ii) Charades-STA [12], (iii) ActivityNet [18], (iv) 100K selected samples from Koala36M [50], and (v) 9K multi-turn dialogues from DSTC10-AVSD [44] enriched with explicit temporal references. The curated dataset comprises 150K QA pairs and is used in DaMO’s training stages 2 and 3. The new benchmark is also released to the research community.
To augment these datasets, we leverage GPT-based prompting to systematically enhance their annotations by generating high-quality, temporally grounded QA pairs. Our contributions are twofold: (1) transforming timestamped event descriptions into natural QA format, and (2) integrating explicit temporal references into multi-turn dialogues. This augmentation provides more precise supervision for training temporal reasoning models.
@@ 7 GPT-augmented Temporal QA Datasets
For datasets with event-level temporal spans, such as Charades-STA [12], we design prompts that instruct GPT to convert timestamped segments into natural questions and corresponding answers. An example is in Fig. 6. This enables learning temporal grounding in a QA context, bridging the gap between annotation spans and linguistic query-answer structures.
Figure 6: GPT-based QA generation from timestamped descriptions.
@@
For multi-turn dialogues with associated temporal segments (e.g., DSTC10-AVSD  [44]), we use GPT to integrate these timestamps into the dialogue naturally. An example is in Fig. 7. The prompt instructs GPT to revise or regenerate responses so that they include explicit temporal references tied to the original annotations, improving temporal traceability and alignment throughout the conversation.
Figure 7: GPT-based temporal enrichment in multi-turn dialogue.

### v3

@@ Abstract
Large Language Models (LLMs) have recently been extended to the video domain, enabling sophisticated video-language understanding. However, existing Video LLMs often exhibit limitations in fine-grained temporal reasoning, restricting their ability to precisely attribute responses to specific video moments, especially under constrained supervision. We introduce DaMO, a data-efficient Video LLM explicitly designed for accurate temporal reasoning and multimodal understanding. At its core, the proposed Temporal-aware Fuseformer employs a hierarchical dual-stream architecture that progressively captures temporal dynamics within each modality and effectively fuses complementary visual and audio information. To further enhance computational efficiency, DaMO integrates a global residual that reduces spatial redundancy while preserving essential semantic details. We train DaMO via a structured four-stage progressive training paradigm, incrementally equipping the model with multimodal alignment, semantic grounding, and temporal reasoning capabilities. This work also contributes multiple datasets augmented from existing ones with LLM-generated temporally grounded QA pairs for tasks requiring temporal supervision. Comprehensive experiments on temporal grounding and video QA benchmarks demonstrate that DaMO consistently surpasses prior methods, particularly in tasks demanding precise temporal alignment and reasoning. Our work establishes a promising direction for data-efficient video-language modeling. Our project is released at https://github.com/yue-723/DaMO
Figure 1: Qualitative comparison on temporal reasoning in video-grounded QA. Given a temporal question grounded in a video clip, DaMO generates a more precise and temporally aligned response than Video-LLaMA [56] and VTimeLLM [14], showcasing superior temporal understanding.
@@ 1 Introduction
To make these trainings possible, this work also contributes multiple temporal QA datasets by enriching existing datasets through prompting with LLM, serving as a valuable asset for temporal reasoning benchmarking.
Extensive experiments on temporal grounding and video dialogue benchmarks confirm DaMO’s strong performance, especially in tasks requiring fine-grained temporal reasoning under limited data and compute. As shown in Fig. 1, DaMO produces more temporally aligned and coherent responses than existing Video LLMs, highlighting the effectiveness of our fusion architecture and progressive training strategy.
@@
  • To preserve global context while significantly reducing computational cost, we devise a global residual.
  • We present a four-stage progressive training paradigm to systematically enhance DaMO’s multimodal alignment and temporal reasoning capabilities. The training is facilitated by several newly curated datasets with LLM-generated temporal QA pairs.
@@ 3 Multimodal Orchestrator DaMO
Fig. 2 illustrates the architecture of DaMO, which is designed to perform multimodal temporal reasoning over audio-visual inputs. The model takes as input an audio-visual stream along with a user prompt, and outputs a response generated by a pretrained language model. To enable the LLM to reason about the audio-visual stream, the audio-visual stream is decomposed into two separate modalities: a silent video stream and an audio stream. Two pretrained visual and audio encoders are employed to separately extract their modality-specific features. These extracted features undergo a spatial path and a temporal path, and then are processed by Q-Former to project multimodal features into tokens that can adapt to the embedding space of the pretrained LLM accompanied by LoRA.
@@ 3.1 Pretrained Video and Audio Encoders
Each frame $V_{i}$ $(i=1,\dots,N)$ and audio segment $A_{j}$ $(j=1,\dots,M)$ is independently encoded by its corresponding encoder. In our implementation, each visual frame is resized to $336\times 336$, and each audio segment length $S=30$ seconds with $R=16$ KHz. The resulting visual and audio features are denoted respectively as:
@@ 3.2 Spatial Pathway
To effectively compress spatial redundancy, we apply adaptive average pooling to $V^{\text{Local}}$ and $A^{\text{Local}}$ separately, reducing both their lengths from $L_{v}-1$ and $L_{a}$ to the same smaller $L^{{}^{\prime}}$. Simultaneously, the global features $V^{\text{Global}}$ and $A^{\text{Global}}$ are each refined through a lightweight feed-forward network (FFN). Finally, we fuse the compressed local representations with the refined global contexts through a residual addition:
@@
Here, $X=V$ or $A$, $V^{\text{Res}}\in\mathbb{R}^{N\times L^{{}^{\prime}}\times D_{v}}$, and $A^{\text{Res}}\in\mathbb{R}^{M\times L^{{}^{\prime}}\times D_{a}}$.
@@ 3.4 Token Projection and LoRA Adaptation
The FUSION queries from the final T-Fuseformer layer thus serves as a compact, temporally grounded representation of the entire video, which is subsequently projected into the LLM’s semantic space using the Q-Former, followed by a linear transformation. These projected multimodal embeddings serve as soft prompts, which are concatenated with user-generated textual queries and fed into the pretrained LLM. To adapt to these FUSION queries, LoRA [13] is applied to the frozen LLM to generate temporally grounded and contextually precise responses.
@@ 3.5 Data-Efficient Progressive Training Paradigm
We employ the VTG objective to condition the frozen LLM, guiding it to generate temporally coherent and contextually accurate responses. To facilitate this semantic alignment, we curate a training corpus comprising approximately 300K QA pairs. This dataset integrates several sources, including general video-based QA data from VideoInstruct-100K [37] and a refined version of AVSD processed by Macaw-LLM [35]. To enrich the temporal grounding component, we additionally incorporate four temporally annotated datasets: QVHighlight [19], Charades-STA [12], ActivityNet [18], and 100K samples selected from Koala36M [50]. Building upon these resources, we leverage LLM-based prompting to regenerate approximately 150K QA pairs with explicit temporal grounding derived from their original segment annotations. The augmentation details are provided in Supplementary.
By combining general QA datasets with the LLM-augmented datasets, we equip DaMO with diverse semantic understanding and fine-grained temporal alignment capabilities, providing robust representation bridging.
@@
The objective in this stage is to adapt DaMO for temporally grounded multi-turn dialogue, enabling it to produce coherent responses while reasoning over temporally structured video features. Building upon the training outcomes of Stage 3, we continue fine-tuning the full pipeline using the VTG objective. This stage leverages a curated 39K dialogue-oriented dataset, comprising general dialogues from VideoChat2 [24] and ActivityNet-based dialogues adapted from VTimeLLM [14]. Additionally, we augment DSTC10-AVSD [44] with LLM-based prompting to inject temporal annotations aligned with its dialogue flows, ensuring it to support training temporally grounded dialogue generation. The details are provided in Supplementary.
@@ 4.2 Zero-Shot Video Retrieval Benchmarks
We first evaluate DaMO’s capability to extract discriminative multimodal representations. This evaluation specifically assesses the effectiveness of Stage 1 training, focusing on DaMO’s multimodal temporal alignment ability without any retrieval-specific fine-tuning. We test the zero-shot video retrieval task on MSR-VTT [53] and MSVD [3]. Following standard protocols, we report Recall@1, Recall@5, and Recall@10 metrics.
@@
As shown in Table 1, DaMO achieves competitive results despite training on considerably less data. On MSR-VTT, DaMO ranks second only to InternVideo2-6B, while on MSVD, it surpasses InternVideo2-6B at Recall@1 and Recall@5, highlighting its strong generalization ability and data efficiency. These results validate the effectiveness of our temporal fusion architecture and multimodal alignment strategy under resource-constrained conditions.
@@ 4.3 Video-LLM Temporal Grounding Benchmarks
As shown in Table 2, DaMO achieves strong overall performance across both benchmarks, significantly outperforming existing Video LLM methods. These outcomes affirm our approach’s superior temporal localization performance, validating the effectiveness of our temporal fusion mechanism and progressive training paradigm.
4.4 Temporal Dialogue Understanding
@@ 5 Conclusions
Additionally, this work contributes several temporal QA datasets that are enhanced from existing video-language datasets via LLM-based prompting. Further dataset information is available in Supplementary. In the future, we plan to further enhance the semantic comprehension and reasoning capabilities of DaMO beyond temporal aspects and incorporate additional modalities.
@@ 6 Additional Qualitative Examples
7 LLM-augmented Temporal QA Datasets
To train DaMO to learn to reason about temporal information, we need to curate a temporal QA dataset without depending on costly large-scale manual annotations. There exist several event-level temporally annotated datasets. While these datasets contain valuable temporal signals, their original annotations are either limited to span-based descriptions without QA structure, or lack fine-grained temporal grounding within dialogues. This restricts their direct applicability to temporal QA tasks, especially for those requiring alignment between natural language queries and specific video moments. Therefore, we construct a LLM-augmented dataset from the following datasets: (i) QVHighlight [19], (ii) Charades-STA [12], (iii) ActivityNet [18], (iv) 100K selected samples from Koala36M [50], and (v) 9K multi-turn dialogues from DSTC10-AVSD [44] enriched with explicit temporal references. The curated dataset comprises 150K QA pairs and is used in DaMO’s training stages 2 and 3. The new benchmark is also released to the research community.
To augment these datasets, we leverage LLM-based prompting to systematically enhance their annotations by generating high-quality, temporally grounded QA pairs. Our contributions are twofold: (1) transforming timestamped event descriptions into natural QA format, and (2) integrating explicit temporal references into multi-turn dialogues. This augmentation provides more precise supervision for training temporal reasoning models.
@@ 7 LLM-augmented Temporal QA Datasets
For datasets with event-level temporal spans, such as Charades-STA [12], we design prompts that instruct LLM to convert timestamped segments into natural questions and corresponding answers. An example is in Fig. 6. This enables learning temporal grounding in a QA context, bridging the gap between annotation spans and linguistic query-answer structures.
Figure 6: LLM-based QA generation from timestamped descriptions.
@@
For multi-turn dialogues with associated temporal segments (e.g., DSTC10-AVSD  [44]), we use LLM to integrate these timestamps into the dialogue naturally. An example is in Fig. 7. The prompt instructs LLM to revise or regenerate responses so that they include explicit temporal references tied to the original annotations, improving temporal traceability and alignment throughout the conversation.
Figure 7: LLM-based temporal enrichment in multi-turn dialogue.